VIF En français | In english
Home | My account | Support | Contacts
Tuesday, December 12 2017 20:15:48 EST
search for on  
This page contains a summary of several connexion tests performed on key components of the network.

Server Status


www.vif.com (HTTP) Good
mail.vif.com (SMTP) Good
ns10.vif.com (DNS) Good
ns20.vif.com (DNS) Good
pop.vif.com (POP3) Good
email.vif.com (Webmail) Good
srs.vif.com (Domain reg.) Good
Extern. 1 Good
Google (HTTP) Good
Yahoo (HTTP) Good
mail.vif.com (IMAP) Good
ftp.vif.com (FTP) Good
Extern. 2 Good
Hotmail (HTTP) Good
ICQ/AOL (5190) Good
Last update on Tue Dec 12 20:15:00 2017


Network status history:

2017-11-02 08:46:54 EDTThe mail server is back online. All our servers are functional.
2017-11-02 08:38:21 EDTOur mail server is experiencing some issues this morning. We are actively working on fixing it. We will keep you posted on the developments.
2017-09-10 13:40:57 EDTThe Intermittent iussue with the VoIP switch was identified and has been fixed. All services are now fully functional
2017-09-10 13:36:06 EDTUPDATE: Our VoIP services are now functional
2017-09-09 19:22:19 EDTWe are currently investigating intermittent call drops on our VoIP switch. Some calls are being dropped under a minute duration. We are actively working to resolve this issue. We apologize for the inconvenience and we will keep you updated as it is resolved.
2017-05-30 11:50:14 EDTAll systems are running smoothly
2017-05-14 19:00:00 EDTIn order to limit the propagation of the current SMBv1 exploits:
The folowing ports are blocked: UDP ports 137-138 and TCP ports 445,137,139
2017-03-10 10:57:48 ESTAll systems are running
2017-02-16 12:33:03 ESTAll our systems are functioning properly.
2017-01-19 15:04:47 ESTOur phone system rebooted. If you have issues restart your VoIP adapter. All systems are functioning properly.

VIF Internet 368 Notre Dame O suite 200 Montreal QC H2Y 1T9 - 514-353-9988 / sans-frais: 888-321-3737